What is the expected market rate of return if Rowdy Corp. has a beta of 0.8, the risk-free rate of return is equal to 2 percent, and the required rate of return for Rowdy is 10%. Assume that Rowdy’s average rate of return over the last 10 years is 9.5%.
Answer #1

As per CAPM,

Required Rate of Return of stock = Risk Free Rate + Beta*[Expected Return on Market-Risk Free Rate]

10 = 2 + 0.8(M-2)

8/0.8 = M-2

10 + 2 = M

Therefore, Expected Return on Market is 12%